# Multi-Modal Framing Analysis of News

### v2

@@ 6 Discussion
Restricting computational analyses of framing to texts only or single labels alone results in missing information that could reshape one’s interpretation of an article. An article can use several frames within the text itself. Further, analysing images is important because it allows for a more complete analysis of media bias, as textual analysis alone may miss the ideological and emotional undertones of visual elements Wessler et al. (2016); Geise et al. (2025). While analyzing images for frames is difficult, as the task is subjective, we find that it is possible reach a meaningful agreement between trained annotators and detect frames with reasonable accuracy across large and diverse datasets.
Our analysis of multimodal framing in the news reveals that more frames are predicted for texts compared to images, that there are distinct lexical patterns associated with text frames vs. image frames, and that frames can be used for detailed topic analysis, including both at a corpus level and at a more fine-grained level by focusing on a single issue. Finally, we show how different news publishers from different ends of the political spectrum use different issue-specific frames to build a narrative, portraying different sides of a story to their readers. The differences we uncover between the use of text and image frames emphasize that it is crucial to take images into account when analyzing framing in the news.
@@ 7 Conclusion
Framing analysis is an important task, with implications for several fields in the social sciences and for understanding society at large. Through our work, we provide a dataset and leverage large vision-language and language models to conduct a more integrative framing analysis, which we hope is useful for future studies in framing analysis in computational and social sciences. This allows for more efficient and replicable analyses of large-scale media corpora. We also urge practitioners working on these topics to look at framing at a broader level and conduct multi-faceted analyses, taking into account both generic and issue-specific frames, visual framing, and portrayal of subjects in their data. Future research should explore human-in-the-loop approaches to ensure both efficiency and interpretive depth. As digital media continues to expand, automated approaches will be crucial for capturing and analyzing the evolving landscape of visual and textual framing across the news ecosystem.
8 Acknowledgments
This work is partially funded by a DFF Sapere Aude research leader grant under grant agreement No 0171-00034B, as well as by the Pioneer Centre for AI, DNRF grant number P1.
@@ 9 Limitations
Despite the advancements in automating multimodal framing analysis, our approach has certain limitations. Firstly, our method assumes that both textual and visual elements contribute meaningfully to framing, yet in some cases, framing might be driven predominantly by one modality. This is akin to a news reader focusing more on one modality while consuming news. The challenge of evaluating the dominance of one modality remain an open problem when text and image signals conflict or reinforce different interpretations. Secondly, the interpretability of automated framing predictions remains limited. While our approach can identify patterns aligned with theoretical framing literature, it does not fully replicate the depth of qualitative human analysis. Ensuring transparency in model decisions and improving explainability remains an important avenue for future research. Finally, our study focuses on a specific set of framing methodologies and datasets, meaning generalizability to other media contexts or platforms requires further investigation. The dynamic nature of framing, influenced by evolving cultural and political landscapes, suggests that models must be continuously updated to remain effective.
@@ Appendix A Source selection
Appendix B Data Filtering
The scraped data carried a lot of noise. There were single line or extremely long articles, images that only depicted logos of news websites, which are not useful for analysis. To remove these, we filtered the articles whose lengths below the 5th percentile and above the 95th percentile, removing the outliers. For images, we similarly removed the files for which the size was above the 95th percentile. There was no clear way to remove the images with logos. To approximate this, we remove all figures with a size of less than 5000 bytes. We also removed articles that were not in English, an information available via news-please library.
Appendix C Frame Descriptions
@@
Appendix D Experimental Details
@@
Pixtral: We used Pixtral-12B-2409 available via hugging face  [^6] https://huggingface.co/mistralai/Pixtral-12B-2409. Similar to Mistral, we used vLLM and set the parameters as: temperature=0.2, max_tokens=1024, dtype="half" and max_model_len=7000. Before processing the images via pixtral, we also resized the images to 512 x 512 for computational efficiency.
@@
Appendix E Article Subject Portrayal
@@
We additionally explore how individual entities in the articles are portrayed. For this analysis, we take the subset where the article and the image are portraying the same main entity. The entities and sentiment are extracted by prompting the LLM and VLM to identify the main subject in the article or the image, as shown in the prompt in 1. Figure 14 shows peculiar differences when contrasting the image portrayal from the portrayal in the articles. A general pattern that can be observed is that entities were portrayed more positively in the images, compared to the text, e.g., Rishi Sunak, Prince Harry, Elon Musk, Joe Biden, Benjamin Netanyahu. For some, there is a negative portrayal overall that is much more explicit in the text, e.g., Donald Trump, Hunter Biden, Vladimir Putin, Rudy Guliani.
Appendix F Frame Classifier Performance
@@
We provide the per label performance of the framing classifier for text on the MFC dataset in Table 4.
Appendix G Additional Plots

### v3

@@ 6 Discussion
Restricting computational framing analysis to text or single labels overlooks important information, as articles often contain multiple frames, and images might convey ideological or emotional undertones that text alone may miss Wessler et al. (2016); Geise et al. (2025). While analysing framing in images is subjective, we find it possible to reach meaningful agreement among trained annotators and detect frames with reasonable accuracy at scale.
Our analysis of multimodal framing in the news reveals that more frames are predicted for texts compared to images, that there are distinct lexical patterns associated with text frames vs. image frames, and that frames can be used for detailed topic analysis, including both at a corpus level and at a more fine-grained level by focusing on a single issue. We show how news publishers across the political spectrum use different issue-specific frames to shape narratives, potentially reinforcing echo chambers and influencing misinformation and societal polarisation. The differences we uncover between the use of text and image frames emphasize that it is crucial to take images into account when analysing framing in the news.
The cause of these different framings in images compared to the text is multi-faceted as they stem from multiple factors, including reliance on copyright-free images, algorithmic bias in image search, editorial intent, individual and institutional biases, and corporate ownership structures. Prior work classifies this as frame-building which shapes how framing appears in news  (De Vreese, 2005). Our method allows for an analysis of the outcome of this process, allowing for a more thorough investigation and analysis at scale.
@@ 7 Conclusion
Framing analysis is an important task for content analysis at large, with implications for several fields in the social sciences, particularly for media bias analysis. In this work, we propose a methodology for leveraging open-weight large (vision-) language models to conduct integrative framing analysis of news articles at scale. While models are highly sensitive to prompting and over-rely on surface level visual features, we show that they can be be prompted to have reasonable accuracy on human-annotated framing datasets. Our approach allows for efficient, customisable, and replicable analyses of images and text in large-scale corpora. We demonstrate the usefulness of our approach by conducting multi-faceted framing analyses, taking into account both generic and issue-specific frames, textual and visual framing, and multiple framings per article. We find stark differences in visual and textual framing of the same topics, as well as differences in framing across publications across the political spectrum. We release a large scale dataset of 500k news articles with model annotations for article text as well as images, which we hope is useful for future studies in framing analysis.
@@ 8 Limitations
Despite the advancements in automating multimodal framing analysis, our approach has certain limitations. Firstly, our method assumes that both textual and visual elements contribute meaningfully to framing, yet in some cases, framing might be driven predominantly by one modality. This is akin to a news reader focusing more on one modality while consuming news. The challenge of evaluating the dominance of one modality remain an open problem when text and image signals conflict or reinforce different interpretations. Similarly, our approach also assumes equivalence across several frame predictions within an article i.e. all predicted frames have the same salience, however, in reality, one frame might be more dominant over another within the text or the image. We chose this design due to method constraints, accurately predicting multiple frames in an article or an image is already an very hard problem, detecting their relative salience is much harder. We ran some initial experiments with naively prompting for predictions in order of saliency but found the outputs to be quite noisy while doing qualitative analysis. Further, the outputs were hard to assess systematically since no such dataset with dominance scores exists for frame prediction. Another limitation is that the interpretability of automated framing predictions remains limited. While our approach can identify patterns aligned with theoretical framing literature, it does not fully replicate the depth of qualitative human analysis. Ensuring transparency in model decisions and improving explainability remains an important avenue for future research. Finally, our study focuses on a specific set of framing methodologies and datasets, meaning generalizability to other media contexts or platforms requires further investigation. The dynamic nature of framing, influenced by evolving cultural and political landscapes, suggests that models must be continuously updated to remain effective.
@@ Appendix A Source selection
Appendix B Extracted Aspects
In Table 3, we provide a list of extracted aspects per modality. For the main subject of the articles, the models were instructed to extract entities central to the text or the image, if there is one. We extract captions from the images and release it as part of the dataset but do not use it for our subsequent framing analysis.
Modality Aspect Description
Text Topic Broad topic of the article
Main subject Subject of the text
Generic Frame One or more of the 15 generic media frames
Issue Frame Inductive, open-ended framing of a topic
Image Caption Caption of the image
Main Subject Subject in the image
Generic Frame One or more of the 15 generic media frames
Table 3: List of aspects extracted per modality.
@@ Appendix B Extracted Aspects
Appendix C Data Filtering
The scraped data carried a lot of noise. There were single line or extremely long articles, images that only depicted logos of news websites, which are not useful for analysis. To remove these, we filtered the articles whose lengths below the 5th percentile and above the 95th percentile, removing the outliers. We also removed articles that were not in English, an information available via news-please library. For images, we similarly removed the files for which the size was above the 95th percentile. There were several images in the dataset that were only logos of news organisations due to the scraper picking the wrong image from the webpage. These only had the logo and no other content, making them irrelevant for our framing analysis. There was no simple way to remove these images, we approximated their identification based on the image size. We distributed the images per their size into bins and plotted 50 randomly sampled images from them. After qualitatively assessing images from each bin, we concluded that 90% of the visualised images below the 5000 bytes threshold mostly constituted of logos only, while it was much lower for higher threshold. Therefore, we remove all figures with a size of less than 5000 bytes.
Appendix D Frame Descriptions
@@
Appendix E Experimental Details
@@
Pixtral: We used Pixtral-12B-2409 available via hugging face  [^6] https://huggingface.co/mistralai/Pixtral-12B-2409. Similar to Mistral, we used vLLM and set the parameters as: temperature=0.2, max_tokens=1024, dtype=‘half’ and max_model_len=7000. Before processing the images via pixtral, we also resized the images to 512 x 512 for computational efficiency.
@@
Appendix F Article Subject Portrayal
@@
We additionally explore how individual entities in the articles are portrayed. For this analysis, we take the subset where the article and the image are portraying the same main entity. The entities and sentiment are extracted by prompting the LLM and VLM to identify the main subject in the article or the image, as shown in the prompt in 1. Figure 14 shows peculiar differences when contrasting the image portrayal from the portrayal in the articles. A general pattern that can be observed is that entities were portrayed more positively in the images, compared to the text, e.g., Rishi Sunak, Prince Harry, Elon Musk, Joe Biden, Benjamin Netanyahu. For some, there is a negative portrayal overall that is much more explicit in the text, e.g., Donald Trump, Hunter Biden, Vladimir Putin, Rudy Giuliani.
Appendix G Frame Classifier Error Analysis
Text
@@ Appendix G Frame Classifier Error Analysis
We provide the per label performance of the framing classifier for text on the MFC dataset in Table 4
We additionally show examples of misclassifications on the dataset in Table 9. Here, while the intersection of the prediction and label sets is null, the explanation of the model for identifying those framings is reasonable and faithful to the article text. To understand patterns in these misclassifications, we analyse labels that the model often gets confused with. Since we’re operating in a multi-label setting, a traditional confusion matrix cannot be applied. A multi-label confusion matrix also treats the labels as binary and not give insight into confusion patterns. To provide insight into these specific confusions, we calculate a mismatch frequency matrix where for each missed ground truth label in an instance (rows in the table below), we mark the erroneous predictions as 1 (columns). We then sum these for the entire dataset, giving us an overview of which labels are often mis-predicted for each missed gold label. For the text frames, the mismatch frequency against MFC is shown in Table 7. The most frequent mismatch is with the political frame, with the model predicting legality, policy or crime framing instead. Culture is another frame often mislabeled, albeit with much lower frequencies.
Image
Similarly, for the image error analysis, we calculate the mismatch frequency matrix as shown in Table 8. As can be seen, most mispredictions involve the None label, with the model over or underpredicting it. Other labels often confused are public opinion and policy prescription, the latter being particularly hard to detect in images even during human annotations.
Examining the  15% of examples where the model’s predicted frames had no overlap with the human-labeled (gold-standard) frames, and we will add this error analysis to the paper. Among these zero-overlap cases, the most common human-assigned label was ‘none’, which accounted for 40.9% of the instances. This indicates that in many such cases, the model over-predicted by assigning frames even when annotators found no meaningful framing. The most frequently missed substantive frame was ‘public opinion’ (11.8%), followed by ‘cultural identity’ (10.8%). Other frames commonly missed included ‘quality of life’, ‘political’, ‘capacity and resources’, and ‘health and safety’, each appearing in approximately 7.5% of the zero-overlap cases. Less frequently missed frames were ‘external regulation and reputation’ (3.2%), and ‘economic’, ‘morality’, and ‘crime and punishment’, each at 1.1We also note that some frames were relatively infrequent in the human-labeled data overall. For instance, ‘fairness and equality’ appeared in only 4.1% of annotated examples, followed by ‘morality’ (2.31%), ‘external regulation and reputation’ (2.31%), ‘economic’ (1.92%), ‘legality, constitutionality and jurisprudence’ (1.92%), and ‘policy prescription and evaluation’ (1.54%). As a result, their low miss rates may reflect either their lower frequency in the dataset or higher accuracy in the model’s predictions.
none economic cap&res morality fairness legality policy crime security health quality_life culture public_op political regulation none 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 economic 5 0 6 198 724 897 970 798 167 178 152 152 408 261 180 cap&res 20 1205 0 95 438 953 1497 380 349 913 854 213 443 140 309 morality 2 207 55 0 340 318 392 282 73 103 164 163 235 150 69 fairness 0 80 20 120 0 177 246 189 37 32 55 88 139 148 29 legality 2 296 153 421 664 0 479 329 182 100 165 285 420 265 100 policy 12 282 126 567 1262 1042 0 649 210 175 308 426 424 478 132 crime 1 204 72 158 498 364 387 0 98 64 80 164 268 349 77 security 1 341 108 117 524 866 809 806 0 60 183 148 397 276 182 health 8 541 194 435 871 1065 1218 1074 507 0 448 300 669 341 194 quality_life 8 639 175 563 1462 713 1181 445 197 237 0 563 892 505 181 culture 15 797 241 576 1036 1189 1766 1179 390 438 634 0 897 563 319 public_op 4 278 77 383 954 951 796 846 219 127 289 364 0 406 142 political 21 1529 406 535 1944 3172 3199 2308 1143 824 1136 569 1243 0 421 regulation 6 401 62 52 220 399 529 205 184 168 221 108 179 79 0
Table 7: Mismatch frequency matrix for the text annotations on the MFC dataset
cap&res crime culture economic fairness health legality morality none policy political public_op quality_life regulation security cap&res 0 3 0 0 0 3 1 1 14 2 0 4 1 0 6 crime 0 0 0 0 0 0 5 0 3 0 0 1 0 0 0 culture 2 0 0 0 4 3 2 3 16 2 0 3 2 0 4 economic 0 1 0 0 0 0 0 0 2 0 0 0 1 0 0 fairness 0 0 0 0 0 0 2 0 1 1 0 0 0 0 0 health 2 4 1 0 0 0 3 3 11 2 0 4 1 0 7 legality 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 morality 0 1 0 0 0 0 0 0 0 0 0 0 0 0 1 none 4 3 7 0 3 4 9 2 0 21 4 6 4 0 8 policy 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 political 0 0 0 0 1 0 3 2 9 9 0 0 0 1 3 public_op 1 0 1 0 0 0 6 0 7 11 1 0 0 1 0 quality_life 3 0 1 0 2 3 2 3 21 1 0 2 0 0 0 regulation 1 0 0 0 0 0 1 1 2 2 0 0 0 0 3 security 0 1 0 0 0 0 2 0 2 0 0 1 0 0 0
Table 8: Mismatch Matrix for Vision annotations across the human annotated dataset
Topic Prediction Label Model reasoning
gun_control [’crime’, ’policy’, ’security’] [’political’, ’health’] The article discusses a gun ban for people on the no-fly list, which is related to crime and punishment and policy prescription and evaluation. Additionally, the context of the discussion is in response to a security threat, making it relevant to the Security and defense frame.
climate_change [’economic’, ’health’, ’quality_life’, ’policy’, ’regulation’] [’cap&res’, ’political’] The article discusses the warming of the oceans over the last 40 years, which has implications for economic costs (e.g., potential damage to marine life and industries), health and safety (e.g., potential impact on human health and marine ecosystems), quality of life (e.g., potential effects on the overall well-being of individuals and communities), policy prescription and evaluation (e.g., the study provides evidence for or against computer models used to predict Earth’s warming), and external regulation and reputation (e.g., international implications of climate change and the role of the U.S. in addressing it).
gun_control [’economic’, ’crime’, ’policy’, ’public_op’, ’culture’] [’regulation’, ’political’, ’quality_life’] The article discusses the financial implications of gun violence (Economic), the legal and enforcement aspects of gun control (Crime and punishment), the proposed policy changes to address the issue (Policy prescription and evaluation), the public opinion and sentiments towards gun control (Public sentiment), and the cultural differences between Japan and America in terms of gun control (Cultural identity).
climate_change [’economic’, ’security’, ’quality_life’] [’cap&res’, ’health’, ’culture’] The article discusses the potential impact of global warming on Antarctica’s ice, which could lead to a rise in sea levels around the world, affecting coastal cities and potentially causing economic, security, and quality of life issues. Additionally, the article mentions the climate change aspect related to the melting of Antarctica’s ice.
Tobacco [’health’, ’legality’, ’fairness’] [’policy’, ’economic’, ’quality_life’] The article discusses the health implications of smoking, the legality of banning smoking in bars and restaurants, and the potential for unequal treatment of smokers compared to non-smokers.
climate_change [’economic’, ’legality’, ’policy’] [’political’, ’regulation’, ’cap&res’] The article discusses the financial implications of adhering to the Kyoto treaty on global warming for the U.S. economy (Economic). The President’s statement that he couldn’t in good faith sign Kyoto because it does not apply to other nations implies a discussion of the treaty’s legality and constitutionality (Legality, constitutionality and jurisprudence). The article also mentions the White House disputing the science behind the conclusion that greenhouse gas emissions contribute to global warming, which suggests a policy prescription and evaluation (Policy prescription and evaluation).
Immigration [’economic’, ’legality’, ’fairness’, ’policy’, ’public_op’] [’quality_life’, ’morality’, ’political’] The article discusses the financial implications of cutting federal assistance to legal immigrants (Economic), the constitutionality of denying assistance to legal residents (Legality, constitutionality and jurisprudence), the balance or distribution of rights and resources for legal immigrants compared to other residents (Fairness and equality), the discussion of specific policies aimed at addressing the issue (Policy prescription and evaluation), and the public sentiment towards allowing elderly immigrants to starve or children to live on the street (Public sentiment).
climate_change [’economic’, ’policy’, ’legality’, ’public_op’, ’political’, ’regulation’] [’culture’, ’cap&res’, ’health’] The article discusses the decision of a TV station manager to not cover global warming stories, citing financial implications (Economic), the evolution of the issue into politics (Policy prescription and evaluation, Political), the perceived lack of conclusive scientific evidence (Legality, constitutionality and jurisprudence), public sentiment towards the issue (Public sentiment), and international implications of the U.S.’s stance on global warming (External regulation and reputation).
Immigration [’morality’, ’political’, ’culture’, ’crime’] [’public_op’, ’fairness’, ’security’] The article discusses the protest against a border patrol group, which is framed as racist and un-American, thus invoking moral and political implications. The protest is organized by a Latino civil rights organization, indicating cultural identity. The group’s activities are framed as a form of policing (crime and punishment), as they plan to patrol the border and ensure immigrants follow certain procedures.
Table 9: Example misclassifications on the MFC dataset along with the model reasoning for the predictions
Appendix H Issue Frame Analysis
As outlined in subsection 4.1, we extract open-ended issue specific frames by provided the model with example issue frames for some issues and instructed to identify and generate appropriate framing of the issue in the article, based on its topic. The generated outputs are limited to few words, but are open-ended as the model is not instructed to choose from a given set of frames but rather generates them. Issue frames are tailored towards fine-grained analysis of specific issues, so they do not repeat a lot. In total, our dataset has over 56k unique issue frames. Some, however, do repeat across topics, we show the top 20 predicted issue frames by frequency, along with the top 3 broad level topics across which they appear in Table 10. As is evident from the table, the overall frequencies are low in the context of the dataset. Most issue-frames are spread across multiple topics rather than being concentrated into one. Some exceptions include Political Crisis, Corruption, Manipulation, and Power Struggle, for which Politics is the major topic across which they appear, which is intuitive. Others like Public Safety Concern are also concentrated under Crime.
Issue Frame Count Top Topics
Humanitarian Crisis 1723 [(’War’, 437), (’Politics’, 284), (’Immigration’, 272)]
Political Crisis 1470 [(’Politics’, 1071), (’Legal’, 294), (’Immigration’, 49)]
Public Health Crisis 1176 [(’Health’, 797), (’Environment’, 155), (’Crime’, 84)]
Political Persecution 1134 [(’Politics’, 606), (’Legal’, 519), (’Crime’, 2)]
Political Corruption 1064 [(’Politics’, 843), (’Legal’, 207), (’Crime’, 12)]
Public Safety Concern 1053 [(’Crime’, 732), (’Legal’, 42), (’Safety’, 37)]
Political Manipulation 931 [(’Politics’, 667), (’Legal’, 179), (’Immigration’, 44)]
Natural Disaster 907 [(’Environment’, 328), (’Weather’, 253), (’Disaster’, 105)]
National Security Threat 824 [(’Politics’, 375), (’War’, 180), (’Immigration’, 63)]
Political Power Struggle 736 [(’Politics’, 649), (’Legal’, 60), (’War’, 10)]
Cultural Celebration 627 [(’Culture’, 285), (’Entertainment’, 178), (’no_topic’, 58)]
Political Scandal 622 [(’Politics’, 446), (’Legal’, 166), (’Crime’, 3)]
Security Threat 604 [(’War’, 244), (’Politics’, 139), (’Immigration’, 80)]
Natural Disaster Threat 589 [(’Environment’, 297), (’Weather’, 112), (’Natural Disasters’, 79)]
Economic Struggle 587 [(’Economy’, 364), (’Business’, 124), (’Politics’, 38)]
Economic Burden 587 [(’Economy’, 152), (’Business’, 128), (’Immigration’, 69)]
Financial Opportunity 578 [(’Finance’, 268), (’Business’, 229), (’Economy’, 48)]
Tragedy 566 [(’Crime’, 217), (’Accident’, 108), (’Transportation’, 46)]
Legal Battle 558 [(’Legal’, 460), (’Politics’, 69), (’Crime’, 11)]
Criminal Threat 557 [(’Crime’, 512), (’Immigration’, 26), (’Legal’, 9)]
Table 10: Most frequent issue-frames across the dataset along with the top 3 article topics they are encountered in
Appendix I Additional Plots